OpenStudy (anonymous):

Find the volume of the square pyramid shown. Round to the nearest tenth if necessary.

OpenStudy (anonymous):



OpenStudy (calculusxy):

The volume of the square pyramid is calculated with the formula: 1/3(b)^2(h) Where b=base and h=height.

OpenStudy (calculusxy):

@Jackilouhoo53

OpenStudy (anonymous):

I know how to calculate it. But how do I find the height? And what would the base be?

OpenStudy (jdoe0001):

hmmm did you see the picture? just use the pythagorean theorem

OpenStudy (anonymous):

a^2 + b^2 = c^2 8^2 + b^2 = 17^2 64 + b^2 = 289 -64 -64 ---------------- b^2 = 225 (square root) 225 =15 So 15 is the height?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

and the square bottom has an area of 16 * 16

OpenStudy (anonymous):

Thank you so much!! (:

OpenStudy (jdoe0001):

yw